if x^2 + y^2 = 4 (x+y-2),then value of (x+y)^2 is



(A) 4    (B) 0    (C) 16  (D) 20

Grade:12

1 Answers

piyush garg
14 years ago

x^2+y^2=4x+4y-8 this represents apoint circle whose radius is zero and centre is (2,2)

(x-2)^2 +(y-2)^2=0 hence it represents apoint x=2 and y=2 therefore (x+y)^2 is (2+2)^2 =16 ans